# Power law $\alpha$-Starobinsky inflation

### v1

@@ A. Background Equations
$S_{E}=\int d^{4}x\,\sqrt{|g|}\left(\frac{1}{2}{\tilde{R}}+\frac{1}{2}\nabla_{\mu}\chi\nabla^{\mu}\chi-V(\chi)\right).$ (14)
@@
$H^{2}$ $=$ $\frac{1}{3}\left(\frac{1}{2}\dot{\chi}^{2}+V(\chi)\right).$ (15)
@@
$\chi^{\prime\prime}+\left(\frac{H^{\prime}}{H}+3\right)\chi^{\prime}+\frac{1}{H^{2}}\frac{dV(\chi)}{d\chi}=0,$ (20)
@@ B. Mukhanov-Sasaki Equations
$ds^{2}=(1+2\chi)dt^{2}+2a(\partial_{i}B-S_{i})dx^{i}dt-a^{2}(1-2\Psi)\delta_{ij}dx^{i}dx^{j}-a^{2}\left(2\partial_{i}\partial_{j}E+2\partial_{(i}F_{j)}+h_{ij}\right)dx^{i}dx^{j},$ (22)
@@
$S_{(2)}=\frac{1}{2}\int\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\left(\frac{du}{d\tau}\right)^{2}-\left(\partial_{i}u\right)^{2}+u^{2}\frac{1}{z}\frac{d^{2}z}{d\tau^{2}}\right]$ (24)
@@
$u_{k}^{\prime\prime}+\left(\frac{H^{\prime}}{H}+1\right)u_{k}^{\prime}+\Biggl{\{}\frac{k^{2}}{a^{2}H^{2}}-\left[2-4\frac{H^{\prime}}{H}\frac{\chi^{\prime\prime}}{\chi^{\prime}}-2\left(\frac{H^{\prime}}{H}\right)^{2}-5\frac{H^{\prime}}{H}-\frac{1}{H^{2}}\frac{d^{2}V}{d\chi^{2}}\right]\Biggr{\}}u_{k}=0.$ (26)
@@ C. Power spectra
$n_{s}$ $=$ $1+\frac{d\ln\mathcal{P_{\mathcal{R}}}}{d\ln k},$ (32)
$n_{t}$ $=$ $\frac{d\ln\mathcal{P}_{t}}{d\ln k},$ (33)
@@
$r=\frac{\mathcal{P}_{t}}{\mathcal{P_{\mathcal{R}}}},$ (34)
@@ IV Observational constraints
To compute the power spectra for scalar perturbations, (30), and tensor perturbations, (31) we modify publicly available code ModeChord, an updated version of ModeCode Mortonson:2010er for the power law $\alpha$-Starobinsky potential (6). ModeCode solves background and perturbations equations, described in the previous section, numerically without using slow-roll approximation for a given inflaton potential. We also vary $N_{pivot}$, the number of e-foldings from the end of inflation the the time when the length scales corresponding to Fourier mode $k_{pivot}$ leave the inflationary horizon, along with the other potential parameters to take into account the general reheating scenario. The primordial power spectra calculated by ModeCode are used in CAMB Lewis:1999bs to compute the angular power spectra for CMB temperature anisotropy and polarization for a given set of cosmological parameters. These theoretically computed angular power spectra are used in CosmoMC Lewis:2002ah , which performs Markov chain Monte-Carlo analysis to put constraints on the parameters of inflaton potential, $N_{pivot}$ and other parameters of $\Lambda$CDM model from various CMB and LSS observations. We have used Planck-2018, BICEP (BK18) BICEP:2021xfz and BAO data to put constraints on the parameter $M$, $\alpha$, $\beta$ of inflaton potential, $N_{pivot}$ and other $\Lambda$CDM parameters. CosmoMC uses Bayesian statistics to compute the posterior probabilities for various parameters, which requires the prior probabilities. We use flat priors for the potential parameters and $N_{pivot}$ given in Table 1, while the priors for other $\Lambda$CDM parameters are used as in Planck:2018vyg . For each parameter the MCMC convergence diagnostic tests is preformed over the four chains using the Gelman and Rubin variance of mean/mean of chain variance R-1 statistics. The marginalized joint probability distributions are plotted using GetDist tool Lewis:2019xzd .
@@
Table 2 shows the observational constraints on the parameters of power law $\alpha$-Starobinsky potential (6) and $N_{pivot}$. obtained from the Planck 2018 observations in combination with BAO and BICEP (BK18) BICEP:2021xfz .
@@
[…]
Fig. 2 illustrates the posterior probability distribution and the confidence regions at $68\%$ and $95\%$ confidence levels for the potential parameter and the derived parameters $(n_{s},r)$. From the Fig. 2, it can be seen that the parameters $(N_{pivot},\,\log_{10}\alpha)$, $(\log_{10}M,\,\log_{10}\alpha)$ and $(\log_{10}M,\,\beta)$ are strongly correlated, while $(\log_{10}M,\,N_{pivot})$, $(\beta,N_{pivot})$ are weekly correlated or almost uncorrelated. However, in case of power law Starobinsky inflation $(\beta,\,N_{pivot})$ and $(\log_{10}M,N_{pivot})$ are strongly correlated Saini:2023uqc , and in case of $\alpha$-Starobinsky inflation $N_{pivot}$ and $\alpha$ are not correlated Saini:2024mun . As depicted in Fig. 2, the derived parameters $n_{s}$ and $r$ are almost uncorrelated, contrary to standard inflationary scenario and $\alpha$-Starobinsky inflation Saini:2024mun , but, similar to power law Starobinsky inflation Saini:2023uqc .
V BAYESIAN MODEL SELECTION
To determine whether the current observations allow deviations from Starobinsky inflation we compute the Bayesian evidences for Starobinsky inflation, power law Starobinsky inflation, $\alpha$-Starobinsky inflation and power law $\alpha$-Starobinsky inflation proposed in this work. In problems of statistical inference in cosmology Bayes’ theorem combines the prior information $\pi(\theta|{\cal M})$ on the parameters $\theta$ of a physical model ${\cal M}$ with the likelihood $\mathcal{L}(y|\theta,{\cal M})$, which represents the distribution of the data points $y$ for a given parameter choice $\theta$, to yield the posterior distribution $p(\theta|y,{\cal M})$,
@@ V BAYESIAN MODEL SELECTION
$p(y|{\cal M})=\int d^{n}\theta\,\mathcal{L}(y|\theta,{\cal M})\pi(\theta|{\cal M})$ (39)
In Bayesian model selection, the posterior probability $p(y|{\cal M})$, which supports a model ${\cal M}$ based on the data $y$, is required to be computed. Bayes’ theorem Trotta:2008qt allows the conditional probability $p(y|{\cal M})$ to be inverted, providing $p({\cal M}|y)$, which is expressed as
$p({\cal M}|y)=\frac{p(y|{\cal M})\pi({\cal M})}{p(y)}.$ (40)
where $\pi({\cal M})$ is the prior probability of the model ${\cal M}$, $p(y|{\cal M})$ is the likelihood of the data $y$ given the model ${\cal M}$, and $p(y)$ is the marginal likelihood or evidence. The normalization is given by the model evidence $p(y)$, which arises from summing over the discrete set of models ${\cal M}_{i}$. Each model ${\cal M}_{i}$ has a corresponding prior probability $\pi({\cal M}_{i})$ and a Bayesian evidence $p(y|{\cal M}_{i})$, as shown in Eq. (39).
$p(y)=\sum_{i}p(y|{\cal M}_{i})\pi({\cal M}_{i})$ (41)
@@ VI Conclusions
In this work we examine a potential formulated by combining the power law and $\alpha$-Starobinsky model, Eq. (6), in the Einstein frame, which contains both the parameters $\alpha$ and $\beta$ to analyze the deviation from Starobinsky inflation in the light of current observation. We obtain the potential for power law $\alpha$-Starobinsky inflation by considering $R^{\beta}$ term in the Einstein action and introducing a parameter $\alpha$ in the conformal transformation via field redefinition motivated by supergravity Cremmer:1978hn ; Cremmer:1982en . We also derive the potential from no-scale supergravity with an underlying noncompact $SU(2,1)/SU(2)\times U(1)$ symmetry. We compute the primordial power spectra numerically using ModeChord 2021ascl by solving background and perturbation equations without using the slow-roll approximation in the Einstein frame. With the help of this we perform MCMC analysis using CosmoMC to put constraints on the parameters $M$, $\alpha$ and $\beta$ of the potential (6) and the number of e-foldings $N_{pivot}$ from the end of inflation to the time when the scales corresponding to $k_{pivot}$ left the Hubble radius during inflation. We find $\log_{10}\alpha=0.37^{+0.82}_{-0.85}$ and $\beta=1.969^{+0.020}_{-0.023}$ at $95\%$ C.L. This indicates that the current CMB and LSS observations allow $2\sigma$ deviation from $\beta=2$. However, the limits on $\log_{10}\alpha$ are larger than the central value indicating that $\alpha=1$ is consistent with the current observations. These results are similar to power law Starobinsky inflation Saini:2023uqc and $\alpha$-Starobinsky inflation Saini:2024mun . We also find the parameter that determines the energy scale of inflation $M=\left(M=3.34^{+2.62}_{-1.73}\right)\times 10^{-5}$ at $95\%$ C. L, and the number of e-folds from the end of inflation to the time when the pivot scale $k_{pivot}$ leaves the inflationary horizon $N_{pivot}=47\pm 10$ at $95\%$ C. L. Interestingly, by performing MCMC analysis, we find that the parameters $\beta$ and $N_{pivot}$ are uncorrelated, which is in contrast to power law Starobinsky inflation Saini:2023uqc where these parameters are strongly correlated. Similarly, the parameters $\log_{10}\alpha$ and $N_{pivot}$ are strongly correlated in contrast to $\alpha$-Starobinsky inflation Saini:2024mun where these are uncorrelated. The parameters $\log_{10}M$ and $N_{pivot}$ are weakly correlated similar to $\alpha$-Starobinsky inflation Saini:2024mun , but, contrary to power law Starobinsky inflation Saini:2023uqc . There is also no correlation between the derived parameters $n_{s}$ and $r$ contrary to $\alpha$-Starobinsky inflation Saini:2024mun , but, similar to power law Starobinsky inflation Saini:2023uqc .
@@
Generalizations of Starobinsky inflation have both observational and phenomenological implications. Recently the Atacama Cosmology Telescope, in combination with Planck 2018, BAO and DESI Y1 data ACT:2025fju , has reported $n_{s}=0.9743\pm 0.0034$, which is slightly higher than the value reported by Planck 2018 observations, i.e., $n_{s}=0.9651\pm 0.004$ Planck:2018jri . This suggests that the Starobinsky model is now disfavored at the $2\sigma$ level ACT:2025tim as it predicts lower values for spectral index. This highlights the need to study various generalizations of Starobinsky inflation. Moreover, the potential for various generalizations of Starobinsky inflation can be incorporated in the framework of supergravity, which can have several implications for particle physics phenomenology Ellis:2019jha ; Ellis:2019opr .

### v2

@@ A. Background equations
$S_{E}=\int d^{4}x\,\sqrt{|g|}\left(\frac{1}{2}{\tilde{R}}+\frac{1}{2}\nabla_{\mu}\chi\nabla^{\mu}\chi-V(\chi)\right),$ (15)
@@
$\quad\quad\quad\quad H^{2}=\frac{1}{3}\left(\frac{1}{2}\dot{\chi}^{2}+V(\chi)\right),$ (16)
and
@@
$\chi^{\prime\prime}+\left(\frac{H^{\prime}}{H}+3\right)\chi^{\prime}+\frac{1}{H^{2}}\frac{dV(\chi)}{d\chi}=0.$ (21)
@@ B. Mukhanov-Sasaki Equations
$ds^{2}=(1+2\chi)dt^{2}+2a(\partial_{i}B-S_{i})dx^{i}dt-a^{2}(1-2\Psi)\delta_{ij}dx^{i}dx^{j}-a^{2}\left(2\partial_{i}\partial_{j}E+2\partial_{(i}F_{j)}+h_{ij}\right)dx^{i}dx^{j}.$ (23)
@@
$S_{(2)}=\frac{1}{2}\int\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\left(\frac{du}{d\tau}\right)^{2}-\left(\partial_{i}u\right)^{2}+u^{2}\frac{1}{z}\frac{d^{2}z}{d\tau^{2}}\right].$ (25)
@@
$u_{k}^{\prime\prime}+\left(\frac{H^{\prime}}{H}+1\right)u_{k}^{\prime}+\Biggl\{\frac{k^{2}}{a^{2}H^{2}}-\left[2-4\frac{H^{\prime}}{H}\frac{\chi^{\prime\prime}}{\chi^{\prime}}-2\left(\frac{H^{\prime}}{H}\right)^{2}-5\frac{H^{\prime}}{H}-\frac{1}{H^{2}}\frac{d^{2}V}{d\chi^{2}}\right]\Biggr\}u_{k}=0,$ (27)
@@ C. Power spectra
The scalar spectral index $n_{s}$ and the tensor spectral index $n_{t}$ are extracted from the numerically obtained power spectra by using their definition as Bassett:2005xm .
$\quad\quad n_{s}=1+\frac{d\ln\mathcal{P_{\mathcal{R}}}}{d\ln k},$ (33)
and
$n_{t}=\frac{d\ln\mathcal{P}_{t}}{d\ln k}.$ (34)
@@
$r=\frac{\mathcal{P}_{t}}{\mathcal{P_{\mathcal{R}}}}.$ (35)
In our analysis both $n_{s}$ and $r$ are derived parameters, and the parameters of the potential (7) $M$, $\alpha$ and $\beta$ are directly constrained from CMB observations.
@@ IV Observational constraints
[…]
@@
Table 2 shows the observational constraints on the parameters of power law $\alpha$-Starobinsky potential (7) and $N_{pivot}$, obtained from the Planck 2018 observations in combination with BAO and BICEP (BK18) BICEP:2021xfz .
@@
Fig. 2 illustrates the posterior probability distribution and the confidence regions at $68\%$ and $95\%$ confidence levels for the potential parameter and the derived parameters $(n_{s},r)$. From the Fig. 2, it can be seen that the parameters $(N_{pivot},\,\log_{10}\alpha)$, $(\log_{10}M,\,\log_{10}\alpha)$ and $(\log_{10}M,\,\beta)$ are strongly correlated, while $(\log_{10}M,\,N_{pivot})$, $(\beta,N_{pivot})$ are weekly correlated or almost uncorrelated. The contours of $(\beta,\log_{10}M)$ are elongated and tilted downward, indicating a negative correlation. As $\log_{10}M$ increases, $\beta$ decreases slightly. For the parameters $(\log_{10}\alpha$, $\log_{10}M)$ contours are stretched and tilted upward, showing a positive correlation. However, in case of power law Starobinsky inflation $(\beta,\,N_{pivot})$ and $(\log_{10}M,N_{pivot})$ are strongly correlated Saini:2023uqc , and in case of $\alpha$-Starobinsky inflation $N_{pivot}$ and $\alpha$ are not correlated Saini:2024mun . Fig. 2 shows the joint $68\%$ and $95\%$ C.L. constraints on scalar spectral index $n_{s}$ and tensor-to-scalar ratio $r$. These are derived parameters, and their constraints are determined by using Eqns. (33) and (35). As depicted in Fig. 2, the derived parameters $n_{s}$ and $r$ are almost uncorrelated, contrary to standard inflationary scenario and $\alpha$-Starobinsky inflation Saini:2024mun , but, similar to power law Starobinsky inflation Saini:2023uqc .
Fig. 3 shows the $r-n_{s}$ predictions for our model with mean values of the potential parameters along with joint $68\%$ and $95\%$ C.L. constraints from ACT, Planck, BAO from DESI and BICEP observations ACT:2025fju ; ACT:2025tim . We choose the parameters $\log_{10}\alpha=0.37$, $\beta=1.969$, $M=10^{-5}$, and vary number of e-foldings $N_{pivot}$ between $40$ to $55$ to determine the dependence of tensor-to-scalar ratio $r$ on $n_{s}$ shown by black line in the figure. It is clear from Fig. 3 that the $r-n_{s}$ predictions of our model agree within $1\sigma$ with the recent ACT observations.
Figure 3: The $r-n_{s}$ predictions for power law $\alpha$-Starobinsky inflation (black line) along with the joint constraints from ACT, Planck, BICEP/Keck and BAO observations ACT:2025tim .
V BAYESIAN MODEL SELECTION
@@
$p(y|{\cal M})=\int d^{n}\theta\,\mathcal{L}(y|\theta,{\cal M})\pi(\theta|{\cal M}).$ (40)
[…]
$p({\cal M}|y)=\frac{p(y|{\cal M})\pi({\cal M})}{p(y)},$ (41)
[…]
$p(y)=\sum_{i}p(y|{\cal M}_{i})\pi({\cal M}_{i}).$ (42)
@@ VI Conclusions
In this work we examine a potential formulated by combining the power law and $\alpha$-Starobinsky model, Eq. (7), in the Einstein frame, which contains both the parameters $\alpha$ and $\beta$ to analyze the deviation from Starobinsky inflation in the light of current observation. We obtain the potential for power law $\alpha$-Starobinsky inflation by considering $R^{\beta}$ term in the Einstein action and introducing a parameter $\alpha$ in the conformal transformation via field redefinition motivated by supergravity Cremmer:1978hn ; Cremmer:1982en . We also derive the potential from no-scale supergravity with an underlying noncompact $SU(2,1)/SU(2)\times U(1)$. We compute the primordial power spectra numerically using ModeChord 2021ascl by solving background and perturbation equations without using the slow-roll approximation in the Einstein frame. With the help of this we perform MCMC analysis using CosmoMC to put constraints on the parameters $M$, $\alpha$ and $\beta$ of the potential (7) and the number of e-foldings $N_{pivot}$ from the end of inflation to the time when the scales corresponding to $k_{pivot}$ left the Hubble radius during inflation. We find $\log_{10}\alpha=0.37^{+0.82}_{-0.85}$ and $\beta=1.969^{+0.020}_{-0.023}$ at $95\%$ C.L. This indicates that the current CMB and LSS observations allow $2\sigma$ deviation from $\beta=2$. However, the limits on $\log_{10}\alpha$ are larger than the central value indicating that $\alpha=1$ is consistent with the current observations. These results are similar to power law Starobinsky inflation Saini:2023uqc and $\alpha$-Starobinsky inflation Saini:2024mun . We also find the parameter that determines the energy scale of inflation $M=\left(M=3.34^{+2.62}_{-1.73}\right)\times 10^{-5}$ at $95\%$ C. L, and the number of e-folds from the end of inflation to the time when the pivot scale $k_{pivot}$ leaves the inflationary horizon $N_{pivot}=47\pm 10$ at $95\%$ C. L. Interestingly, by performing MCMC analysis, we find that the parameters $\beta$ and $N_{pivot}$ are uncorrelated, which is in contrast to power law Starobinsky inflation Saini:2023uqc where these parameters are strongly correlated. Similarly, the parameters $\log_{10}\alpha$ and $N_{pivot}$ are strongly correlated in contrast to $\alpha$-Starobinsky inflation Saini:2024mun where these are uncorrelated. The parameters $\log_{10}M$ and $N_{pivot}$ are weakly correlated similar to $\alpha$-Starobinsky inflation Saini:2024mun , but, contrary to power law Starobinsky inflation Saini:2023uqc . There is also no correlation between the derived parameters $n_{s}$ and $r$ contrary to $\alpha$-Starobinsky inflation Saini:2024mun , but, similar to power law Starobinsky inflation Saini:2023uqc .
Recently the Atacama Cosmology Telescope, in combination with Planck 2018, BAO data from DESI Y1 ACT:2025fju , has reported $n_{s}=0.9743\pm 0.0034$, which is slightly higher than the value reported by Planck 2018 observations, i.e., $n_{s}=0.9651\pm 0.004$ Planck:2018jri . This suggests that the Starobinsky model is now disfavored at the $2\sigma$ level ACT:2025tim as it predicts lower values for spectral index. We also obtain $r-n_{s}$ predictions for our model for the mean values of the potential parameters $\alpha$ and $\beta$ and varying $N_{pivot}$ between $40$ to $55$. We find that these predictions lie with $68\%$ C.L. of joint constraints from ACT, Planck-2018, BICEP and BAO data from DESI Y1 ACT:2025fju ; ACT:2025tim .
@@
Generalizations of Starobinsky inflation have both theoretical and phenomenological implications. As the original Starobinsky model is now disfavored at the $2\sigma$ level by the recent ACT observations ACT:2025tim , it highlights the need to study various generalizations of Starobinsky inflation. Moreover, the potential for various generalizations of Starobinsky inflation can be incorporated in the framework of supergravity, which can have several implications for particle physics phenomenology Ellis:2016ipm ; Ellis:2017jcp ; Ellis:2019jha ; Ellis:2019opr .